OpenStudy (anonymous):

What is the equation of the line in slope-intercept form? the line perpendicular to y =1/3x+5 through (2, 1) A. y = 1/3x + 7 B. y = -3x + 7 C. y = 3x + 7 D. y = -1/3x + 7

OpenStudy (angela210793):

Two lines are perpendicular when the product of their slopes equals -1 so slope1*slope2=-1 y=ax+b, a=slope and b=y-intercept In your problem you have slope1=1/3 so slope2=-1/(1/3)=-3 the equation of the line when you're given the slpoe and a point is y-yo=slope(x-x0) where (x0;y0) are the coordinates of the point given so y-1=-3(x-2) y-1=-3x+6 y=-3x+7

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

B

OpenStudy (anonymous):

So u guys are saying the answer is B

OpenStudy (angela210793):

yes B

OpenStudy (anonymous):

Thank you so much u guys